OpenStudy (anonymous):

K is the midpoint of line AB . If line AK= 3x - 4 and line KB =1/2x+6 , find the length of line AB

OpenStudy (anonymous):

there is a line ab k bisects it this means that ak congruent kb

OpenStudy (anonymous):

i think you can figure out the rest right?

OpenStudy (anonymous):

not really. im stupid:/

OpenStudy (anonymous):

no you are not

OpenStudy (anonymous):

what is the value of ak

OpenStudy (anonymous):

3x-4

OpenStudy (anonymous):

and value of bk

OpenStudy (anonymous):

then put that into an equation

OpenStudy (anonymous):

whats bk

OpenStudy (anonymous):

is ab equal to 4

OpenStudy (anonymous):

KB bk kinda interchange able

OpenStudy (anonymous):

at this point

OpenStudy (anonymous):

is AB=4

OpenStudy (anonymous):

how did you get that ?

OpenStudy (anonymous):

you were right the first time

OpenStudy (anonymous):

it is the given that you typed

OpenStudy (anonymous):

K is the midpoint of line AB . If line AK= 3x - 4 and line KB =1/2x+6 , find the length of line AB

OpenStudy (anonymous):

?

OpenStudy (anonymous):

If line AK= 3x - 4 and line KB =1/2x+6

OpenStudy (anonymous):

i know i solved

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i got 4

OpenStudy (anonymous):

so ya know how it works right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

can you help me on some other questions?